OpenStudy (ram231):

Who understands what log is in math? I have never learned it and it is on my ACT Prep assignment

OpenStudy (ram231):

@Whitemonsterbunny17 @Elsa213 @tatumlee

zepdrix (zepdrix):

The logarithm? It's an operation which is the inverse of the exponential operation. Example:\[\large\rm 2^3=8\]This exponential operation is a base of 2 raised to the 3rd power. And the output is 8. \[\large\rm \log_28=3\]This log operation is a base of 2 and it takes the number 8 and spit outs a 3. It's like it's giving us the exponent as a result. The log is good for dealing with really really large scale numbers. It's how they measure earthquakes and other interesting stuff :)

zepdrix (zepdrix):

Log has some weird properties:\[\large\rm \log(ab)=\log(a)+\log(b)\]\[\large\rm \log\left(\frac{a}{b}\right)=\log(a)-\log(b)\]\[\large\rm \log(a^b)=b\cdot \log(a)\]

OpenStudy (ram231):

then what does \[\log_{10} A/B\]

zepdrix (zepdrix):

If log is showing up on your ACT, they probably want you to apply these three rules some how.

OpenStudy (ram231):

I dont understand

zepdrix (zepdrix):

Is that the entire question? Do they want you to "expand" it?

OpenStudy (ram231):

thats the entire question with options that come with it.

zepdrix (zepdrix):

can you list at least one of the options so I have an idea of what they want? c:

OpenStudy (ram231):

these are the options

zepdrix (zepdrix):

Well did you notice the 3 rules I listed for log a moment ago? :) It looks like that middle one \(\large\rm \log\left(\frac{a}{b}\right)=\log(a)-\log(b)\) is useful for our problem here.

zepdrix (zepdrix):

Replace a with A, and b with B, and give the log a base of 10,\[\large\rm \log_{10}\left(\frac{A}{B}\right)=\log_{10}(A)-\log_{10}(B)\]

zepdrix (zepdrix):

If you haven't done any work with logs, then yes, these are going to be confusing. You just need to try and follow how these three rules work :)

OpenStudy (ram231):

I havent learned any of the cos, sig, or the logs terms. Im not that smart Dx

OpenStudy (ram231):

wait. I don't see an option of what you gave me for the problem.... Dx

zepdrix (zepdrix):

ACT is a good test :) Try to do well on it. You don't want to rely on SAT, that one is a real piece of garbage. You can at least take educated guesses on some of these ^^ hehe

zepdrix (zepdrix):

You don't? It's the option below the one you highlighted. Keep in mind that \(\large\rm \log_{10}(A)-\log_{10}(B)\quad=\log_{10}A-\log_{10}B\) The brackets don't matter.

OpenStudy (ram231):

oh okay. thats what was confusing. thanks! Be my tutor hahaha (:

zepdrix (zepdrix):

XD

OpenStudy (ram231):

what? Ill need it! hahahaa